# Operations & Supply Chain Management: Principles and Practice

### v1

@@ 4 OM-SCM intersection
Ethical sustainability is a concept to identify an organisation’s ability to understand and act on sustainable values and principles within the legal and moral norms of the given context. In contrast to other dimensions (economy, environment, and society) of sustainability, the ethical dimension has often been left out of the debate. Compliance with business ethics within the sustainability context improves brand value and reputation, resulting in good organisational performance and long-term financial gains (Flores-Hernández et al., 2020). If the taxonomy of ethical sustainability digs further, then ’ethics’ can be defined as a practical philosophy that teaches us to practice certain habits and customs that lead to a better life. Within the context of sustainability, however, several researchers (de Bakker et al., 2019; Hockerts and Searcy, 2023) define business ethics as the concept of evaluating moral ‘right’ and ‘wrong’. By understanding this morality, organisational leaders can make ethical decisions, and such ethical behaviour can improve customer satisfaction, shareholder loyalty, and employee satisfaction. Compared to sustainability, ethics in business is an older context. The earliest publication of business ethics can be seen from the early 1960s by (Baumhart, 1961), when the relationship between ethics and the values of business leaders was investigated. Several studies (Donaldson and Dunfee, 1994; Reidenbach and Robin, 1990; Kolk, 2016; Schneider and Murray, 2024; Rydenfelt, 2024) examine different scenarios of business ethics, and the role of sustainability began to gain momentum about thirty years after the Brundtland report was published in 1987 (Khalili et al., 2015; World Commission on Environment and Development, 1987).
@@ Conclusion
Studies on the transition from global to local healthcare supply chains (HSCs) suggest that local procurement is vital in preparing for future disruptions (Harland et al., 2021). For example, in response to the United Kingdom’s (UK) exit from the European Union (BREXIT), the majority of pharmaceutical companies planned to relocate production sites from the UK to EU states and distribution centres from the EU to the UK (Moradlou et al., 2021). Post-pandemic, similar strategies are likely to be adopted more frequently to shorten HSCs and enhance responsiveness and responsiveness (Duong et al., 2024; Finkenstadt and Handfield, 2021; Xu et al., 2020).
@@
For a comprehensive and up-to-date review of the role of operations and supply chains in healthcare, readers are encouraged to consult recent literature. Liu et al. (2024c) present a framework for patient engagement in remote consultations; Moons et al. (2018) analyse of key performance indicators in hospital logistics, Mišić and Perakis (2020) explore data analytics adoption in healthcare; and Machta et al. (2019) study the integration between healthcare providers and hospital systems. Additionally, reviews of empirical (Kc et al., 2020) and modelling studies in HSCM (Ali and Kannan, 2022; Keskinocak and Savva, 2020) offer valuable insights into the current landscape. For a broader strategic perspective, Schneller et al. (2023) provide a comprehensive overview of the issues facing HSCM and its role in healthcare strategy. These works are essential for understanding the modern requirements of HSCM, offering valuable case studies and insights into the complexities of contemporary healthcare supply chains.
@@
A typical supply chain of activities starts with assessing and publishing needs and budgeting (Keshvari Fard et al., 2022), followed by sourcing and procurement, transport, warehousing and distribution. Nowadays, humanitarian supply chains also include waste management. Some of these activities take place locally aiming to increase localisation (Frennesson et al., 2022), while others internationally. Which activities should take place locally and which internationally depends significantly on the available local resources: very often humanitarian organisations operate in countries where key infrastructures such as water, transport, and power are already struggling to meet in-country needs and placing further demand on these systems could give rise to tensions between humanitarian organisations and local communities (Duong et al., 2024). Humanitarian organisations have an increasing understanding of such conditions, shown for example by their decision to implement cash-based assistance or not. The supply chain often uses a push system immediately after the disaster has occurred and changes to pull with more time and better information on the needs.
@@
Humanitarian operations often take place in resource austere settings with limited access to power, clean water, poor transport links, disrupted supply chains, limited cold chain as well as political volatility. Country access, the import and resupply of key equipment and medical consumables, security and communication issues and local environmental conditions all have an impact. The experiences in 2004/2005 clearly demonstrated the need for logistics preparedness (Jahre et al., 2016; Stumpf et al., 2023) and coordination, collaboration and co-operation (Balcik et al., 2010; Jahre and Jensen, 2010; Kovács and Falagara Sigala, 2021) to increase operational efficiency and effectiveness by closing gaps, reducing duplications, and managing resources better. Substantial efforts by the organisations followed. Until 2015 the most common strategy was strategic stock of relief items, but also other ‘stock’ including vehicle fleets (Pedraza Martinez et al., 2011), emergency funds and rosters of trained human resources (Jahre, 2017). The use of postponement through non-earmarking of items, prepositioning of semi-finished goods, centralisation and outsourcing, collaboration (Altay and Pal, 2014; Ergun et al., 2014) as well as use of flexible transportation and supply base have also been common. Providers must navigate a range of extremes, e.g. temperature, humidity, rainfall, etc., all of which can have a profound effect on the ability to effectively deliver humanitarian operations. Humanitarian organisations should always consider whether collaboration is appropriate given the need to maintain integrity and independence, and depending on aspects like interoperability of technologies and systems, the use and implementation of different regulations and standards, and information asymmetry (Phillips et al., 2023). They should also consider more collaboration with the commercial sector: temporary supply chains dismantled once the disruption has been addressed are particularly important during humanitarian events, ensuring the delivery of critical supplies over a limited timeframe (Müller et al., 2023; Duong et al., 2024).
@@
Different foci: Research on humanitarian operations has traditionally focused much on response and preparedness. In the future we should see more on mitigation and recovery. Related to this we will see more focus on sustainability, i.e., going beyond efficiency and effectiveness in the traditional sense. Further, research has focused a lot on cost and response time, while recently more metrics like equity (Breugem et al., 2024) and pain (Holguín-Veras et al., 2013; Wang et al., 2019) are considered important.
@@
Before Purchase: This stage focuses on how consumers acquire product information, become aware of a brand, and decide to purchase. Traditionally, brick-and-mortar stores were the primary points of purchase. With omnichannel retail, these stores now also serve as catalogues where consumers can acquire information by seeing, touching, and comparing products. To capitalise on this, some retailers are opening showrooms that do not carry inventory. Showrooms are associated with higher demand (Bell et al., 2018), fewer returns (Bell et al., 2020), improved operational efficiency (Bell et al., 2018), and lower sensitivity to fulfilment lead times (Lim et al., 2024a). In addition to brick-and-mortar stores and showrooms, consumers use online brands as an information acquisition channel (Gallino and Moreno, 2014), leading to research on online retail design (Shi et al., 2013).
@@ Methods
Retail research employs a wide variety of methods, which we categorise as analytical, empirical, experimental, and numerical. Among analytical models, the traditional optimisation approach is common to provide a decision model to mathematically optimise retail variables, including assortment (Ulu et al., 2012), order replenishment (Balakrishnan et al., 2004), and phantom stockouts (Chen, 2021). Theoretical models are also common. These stylised models are used to model pricing and consumer welfare (Gui et al., 2019), assortment (Heese and Martínez-de Albéniz, 2018), showroom network design (Lim et al., 2024a), and fulfilment (Jin et al., 2018).
@@
Lab and online experiments are widespread given their ease of implementation and low costs. They are used in research on consumer preferences related to sustainability (Hampl and Loock, 2013), information precision (Lim et al., 2024a), and store environment (Zhang et al., 2023b). This experiment type is common in behavioural research, employing decision-making theories (Liang et al., 2024). Field studies are also conducted to analyse retail concepts such as store environment (Zhang et al., 2023b) and fitting rooms (Lee et al., 2021).
@@
Future research can explore last-mile delivery innovations, which are essential to meet growing consumer demand in the retail sector (Lim, 2023). This segment of the supply chain will be stress-tested in coming years. The integration of artificial intelligence into various retailing decisions, such as inventory or pricing, can also be investigated. Furthermore, the extant literature typically employs discrete choice models, such as the multinomial logit, where consumers are assumed to select a single item, or the multivariate logit, where consumers choose multiple items, each in a unit quantity. A more realistic approach is to consider basket-level choices, including varying quantities of multiple items (see Lim et al., 2024b, for an application). We hope these discussions inspire further studies and method contributions to the retail operations literature.

### v2

@@ 4 OM-SCM intersection
Ethical sustainability is a concept to identify an organisation’s ability to understand and act on sustainable values and principles within the legal and moral norms of the given context. In contrast to other dimensions (economy, environment, and society) of sustainability, the ethical dimension has often been left out of the debate. Compliance with business ethics within the sustainability context improves brand value and reputation, resulting in good organisational performance and long-term financial gains (Flores-Hernández et al., 2020). If the taxonomy of ethical sustainability digs further, then ’ethics’ can be defined as a practical philosophy that teaches us to practice certain habits and customs that lead to a better life. Within the context of sustainability, however, several researchers (de Bakker et al., 2019; Hockerts and Searcy, 2023) define business ethics as the concept of evaluating moral ‘right’ and ‘wrong’. By understanding this morality, organisational leaders can make ethical decisions, and such ethical behaviour can improve customer satisfaction, shareholder loyalty, and employee satisfaction. Compared to sustainability, ethics in business is an older context. The earliest publication of business ethics can be seen from the early 1960s by (Baumhart, 1961), when the relationship between ethics and the values of business leaders was investigated. Several studies (Donaldson and Dunfee, 1994; Reidenbach and Robin, 1990; Kolk, 2016; Schneider and Murray, 2024; Rydenfelt, 2025) examine different scenarios of business ethics, and the role of sustainability began to gain momentum about thirty years after the Brundtland report was published in 1987 (Khalili et al., 2015; World Commission on Environment and Development, 1987).
@@ Conclusion
Studies on the transition from global to local healthcare supply chains (HSCs) suggest that local procurement is vital in preparing for future disruptions (Harland et al., 2021). For example, in response to the United Kingdom’s (UK) exit from the European Union (BREXIT), the majority of pharmaceutical companies planned to relocate production sites from the UK to EU states and distribution centres from the EU to the UK (Moradlou et al., 2021). Post-pandemic, similar strategies are likely to be adopted more frequently to shorten HSCs and enhance responsiveness and responsiveness (Duong et al., 2025; Finkenstadt and Handfield, 2021; Xu et al., 2020).
@@
In addition to SCR, localisation and AI, existing studies on healthcare applications over the last decade have covered several themes from the operations and supply chain management perspective. For example, business analytics (Terekhov et al., 2024), capacity strain (Kim et al., 2024), decision making (van Rijn et al., 2024), inventory management (Yang et al., 2025), quality of care (Xu et al., 2025) and sustainability (Berenguer et al., 2024). Furthermore, studies on this topic are mostly empirical (Kc et al., 2020) and modelling works (Ali and Kannan, 2022; Keskinocak and Savva, 2020). For a comprehensive and up-to-date review of the role of operations and supply chains in healthcare, readers are encouraged to consult recent literature. Liu et al. (2024c) present a framework for patient engagement in remote consultations; Moons et al. (2018) analyse key performance indicators in hospital logistics; and for a broader strategic perspective, Schneller et al. (2023) provide a comprehensive overview of the issues facing HSCM and its role in healthcare strategy. These works are essential for understanding the modern requirements of HSCM, offering valuable case studies and insights into the complexities of contemporary healthcare supply chains. Moreover, they offer some suggestions for future studies. For example, future studies could adopt different methods or mixed methods to explore mechanisms that affect the effectiveness of healthcare operations management (Kc et al., 2020). Also, to increase the research impact in practice, closer collaboration between researchers and practitioners is vital for accessing actual and timely data (Keskinocak and Savva, 2020). Moreover, we call for more interdisciplinary studies, combining healthcare management with fields such as public sector, engineering and pharmaceutical manufacturing.
@@
A typical supply chain of activities starts with assessing and publishing needs and budgeting (Keshvari Fard et al., 2022), followed by sourcing and procurement, transport, warehousing and distribution. Nowadays, humanitarian supply chains also include waste management. Some of these activities take place locally aiming to increase localisation (Frennesson et al., 2022), while others internationally. Which activities should take place locally and which internationally depends significantly on the available local resources: very often humanitarian organisations operate in countries where key infrastructures such as water, transport, and power are already struggling to meet in-country needs and placing further demand on these systems could give rise to tensions between humanitarian organisations and local communities (Duong et al., 2025). Humanitarian organisations have an increasing understanding of such conditions, shown for example by their decision to implement cash-based assistance or not. The supply chain often uses a push system immediately after the disaster has occurred and changes to pull with more time and better information on the needs.
@@
Humanitarian operations often take place in resource austere settings with limited access to power, clean water, poor transport links, disrupted supply chains, limited cold chain as well as political volatility. Country access, the import and resupply of key equipment and medical consumables, security and communication issues and local environmental conditions all have an impact. The experiences in 2004/2005 clearly demonstrated the need for logistics preparedness (Jahre et al., 2016; Stumpf et al., 2023) and coordination, collaboration and co-operation (Balcik et al., 2010; Jahre and Jensen, 2010; Kovács and Falagara Sigala, 2021) to increase operational efficiency and effectiveness by closing gaps, reducing duplications, and managing resources better. Substantial efforts by the organisations followed. Until 2015 the most common strategy was strategic stock of relief items, but also other ‘stock’ including vehicle fleets (Pedraza Martinez et al., 2011), emergency funds and rosters of trained human resources (Jahre, 2017). The use of postponement through non-earmarking of items, prepositioning of semi-finished goods, centralisation and outsourcing, collaboration (Altay and Pal, 2014; Ergun et al., 2014) as well as use of flexible transportation and supply base have also been common. Providers must navigate a range of extremes, e.g. temperature, humidity, rainfall, etc., all of which can have a profound effect on the ability to effectively deliver humanitarian operations. Humanitarian organisations should always consider whether collaboration is appropriate given the need to maintain integrity and independence, and depending on aspects like interoperability of technologies and systems, the use and implementation of different regulations and standards, and information asymmetry (Phillips et al., 2023). They should also consider more collaboration with the commercial sector: temporary supply chains dismantled once the disruption has been addressed are particularly important during humanitarian events, ensuring the delivery of critical supplies over a limited timeframe (Müller et al., 2023; Duong et al., 2025).
@@
Different foci: Research on humanitarian operations has traditionally focused much on response and preparedness. In the future we should see more on mitigation and recovery. Related to this we will see more focus on sustainability, i.e., going beyond efficiency and effectiveness in the traditional sense. Further, research has focused a lot on cost and response time, while recently more metrics like equity (Breugem et al., 2025) and pain (Holguín-Veras et al., 2013; Wang et al., 2019) are considered important.
@@
Before Purchase: This stage focuses on how consumers acquire product information, become aware of a brand, and decide to purchase. Traditionally, brick-and-mortar stores were the primary points of purchase. With omnichannel retail, these stores now also serve as catalogues where consumers can acquire information by seeing, touching, and comparing products. To capitalise on this, some retailers are opening showrooms that do not carry inventory. Showrooms are associated with higher demand (Bell et al., 2018), fewer returns (Bell et al., 2020), improved operational efficiency (Bell et al., 2018), and lower sensitivity to fulfilment lead times (Lim et al., 2024). In addition to brick-and-mortar stores and showrooms, consumers use online brands as an information acquisition channel (Gallino and Moreno, 2014), leading to research on online retail design (Shi et al., 2013).
@@ Methods
Retail research employs a wide variety of methods, which we categorise as analytical, empirical, experimental, and numerical. Among analytical models, the traditional optimisation approach is common to provide a decision model to mathematically optimise retail variables, including assortment (Ulu et al., 2012), order replenishment (Balakrishnan et al., 2004), and phantom stockouts (Chen, 2021). Theoretical models are also common. These stylised models are used to model pricing and consumer welfare (Gui et al., 2019), assortment (Heese and Martínez-de Albéniz, 2018), showroom network design (Lim et al., 2024), and fulfilment (Jin et al., 2018).
@@
Lab and online experiments are widespread given their ease of implementation and low costs. They are used in research on consumer preferences related to sustainability (Hampl and Loock, 2013), information precision (Lim et al., 2024), and store environment (Zhang et al., 2023b). This experiment type is common in behavioural research, employing decision-making theories (Liang et al., 2024). Field studies are also conducted to analyse retail concepts such as store environment (Zhang et al., 2023b) and fitting rooms (Lee et al., 2021).
@@
Future research can explore last-mile delivery innovations, which are essential to meet growing consumer demand in the retail sector (Lim, 2023). This segment of the supply chain will be stress-tested in coming years. The integration of artificial intelligence into various retailing decisions, such as inventory or pricing, can also be investigated. Furthermore, the extant literature typically employs discrete choice models, such as the multinomial logit, where consumers are assumed to select a single item, or the multivariate logit, where consumers choose multiple items, each in a unit quantity. A more realistic approach is to consider basket-level choices, including varying quantities of multiple items (see Lim et al., 2025, for an application). We hope these discussions inspire further studies and method contributions to the retail operations literature.
@@
Data availability statement
No data have been used in this article.